### v1

# Q-Insight: Understanding Image Quality via Visual Reinforcement Learning

Abstract
Image quality assessment (IQA) focuses on the perceptual visual quality of images, playing a crucial role in downstream tasks such as image reconstruction, compression, and generation. The rapid advancement of multi-modal large language models (MLLMs) has significantly broadened the scope of IQA, moving toward comprehensive image quality understanding that incorporates content analysis, degradation perception, and comparison reasoning beyond mere numerical scoring. Previous MLLM-based methods typically either generate numerical scores lacking interpretability or heavily rely on supervised fine-tuning (SFT) using large-scale annotated datasets to provide descriptive assessments, limiting their flexibility and applicability. In this paper, we propose Q-Insight, a reinforcement learning-based model built upon group relative policy optimization (GRPO), which demonstrates strong visual reasoning capability for image quality understanding while requiring only a limited amount of rating scores and degradation labels. By jointly optimizing score regression and degradation perception tasks with carefully designed reward functions, our approach effectively exploits their mutual benefits for enhanced performance. Extensive experiments demonstrate that Q-Insight substantially outperforms existing state-of-the-art methods in both score regression and degradation perception tasks, while exhibiting impressive zero-shot generalization to comparison reasoning tasks. Code will be available at https://github.com/lwq20020127/Q-Insight.
11footnotetext: Project Lead. 🖂: Corresponding authors, zhaoshijie.0526@bytedance.com, zhangjian.sz@pku.edu.cn.
1 Introduction
Image quality assessment (IQA) is a fundamental task in computer vision, critical for optimizing algorithms, enhancing user experiences, and verifying content authenticity across diverse domains, such as image processing [38, 21, 10, 51] and AI-generated content (AIGC) [35, 7]. Traditional IQA methods rely heavily on hand-crafted metrics, either through reference-based comparisons [47] or statistical measures of natural image properties [62]. However, these approaches typically focus on local image characteristics and fail to comprehensively capture global visual quality, limiting their reliability in complex real-world scenarios. More recently, deep learning-based IQA models [44, 18] have emerged, utilizing neural networks to learn hierarchical image representations. Nevertheless, these methods struggle to face significant challenges in out-of-distribution (OOD) generalization.
With recent advances in multi-modal large language models (MLLMs)[22, 1, 45], researchers have begun to leverage these models’ extensive world knowledge and perceptual abilities to enhance IQA performance and broaden its applicability[51, 58, 57, 56, 49]. Existing MLLM-based IQA methods generally fall into two categories: score-based methods, such as Q-Align [51] and DeQA-Score [56], and description-based methods, exemplified by DepictQA [58] and DepictQA-Wild [57]. Score-based methods transform discrete tokens into continuous quality scores, thereby improving adaptability but typically sacrificing interpretability and neglecting MLLMs’ intrinsic reasoning and descriptive capabilities. Meanwhile, simply regressing a quality score may not be meaningful in certain scenarios, as image quality scores are subjective, inherently biased, and lack uniform standards across different datasets and content types. For example, when evaluating AIGC-generated data, unusual visual effects and vibrant colors often imply better quality. However, for evaluating super-resolution results, these same features are often considered too painterly, losing the image’s authenticity and fidelity. Conversely, description-based methods produce detailed textual explanations of image degradations and comparative assessments, ensuring interpretability yet heavily depending on extensive textual depiction for supervised fine-tuning. Moreover, these models cannot output precise scores, making them unsuitable when used as loss functions or for performing accurate ranking of image quality. Consequently, integrating numerical scoring and descriptive reasoning into a unified, interpretable MLLM-based IQA framework remains an essential yet unresolved challenge.
Figure 1: We present the PLCC comparison between proposed Q-Insight and other IQA metrics on the left and list the applications of our Q-Insight on the right. Q-Insight shows significant performance on the out-of-domain datasets (e.g. CSIQ) compared to methods like DeQA, while simultaneously achieving score regression, degradation perception, and zero-shot image comparison reasoning tasks.
In this paper, we move towards a comprehensive understanding of image quality by addressing tasks such as image description, aesthetic and compositional evaluation, degradation perception, and comparative reasoning across images. Rather than teaching the large model “How to score images”, we aim to inspire it “How to reason deeply and formulate insightful perspectives on image quality metrics during scoring”. To this end, we resort to Group Relative Policy Optimization (GRPO) [37], a reinforcement learning framework inspired by DeepSeek-R1 [14]. GRPO has recently shown to be highly effective in large language models (LLMs). It uses heuristic reward signals to efficiently guide LLMs in uncovering their intrinsic reasoning capabilities, removing extensive reliance on annotated reasoning chains or additional value models. Recently, researchers have also successfully adapted GRPO to vision-language tasks, including few-shot object detection, reasoning grounding [24], and medical analysis [33]. In the context of image quality understanding, the introduction of GRPO provides at least three distinct advantages: (1) no reliance on massive textual training data, (2) strong generalization to OOD evaluated images, and (3) high diversity in supporting multiple tasks. These benefits align well with our goal of developing a generalized image quality understanding agent.
Specifically, we design Q-Insight upon the GRPO framework. In our Q-Insight, we jointly optimize score regression and degradation perception tasks, and carefully design three reward functions: a verifiable score reward for the score regression task, and degradation classification and intensity perception rewards for the degradation perception task. Consequently, Q-Insight effectively exhibits robust reasoning performance in image quality assessment tasks using only limited Mean Opinion Scores (MOS) and degradation labels. As shown in Fig. 1, our Q-Insight shows exceptional improvements on OOD datasets, while demonstrating comprehensive quality perception and sophisticated reasoning abilities, such as recognizing that a slightly out-of-focus background, which might conventionally be perceived as undesirable, actually helps emphasize the main subject of the image. Our empirical investigation reveals that: (1) training solely with score labels results in poor perception of image detail degradations (e.g., JPEG compression), while jointly training with the degradation perception task significantly enhances the model’s sensitivity to such degradations, and (2) score regression and degradation perception tasks are mutually beneficial. Extensive experiments across score regression and degradation perception tasks demonstrate that Q-Insight consistently outperforms existing model-based IQA metrics as well as SFT-driven large language models. Moreover, it exhibits impressive zero-shot generalization on unseen tasks, such as image comparison reasoning, highlighting the robustness and versatility of our method. In summary, our contributions are:
❑ (1) We propose Q-Insight, the first reasoning-style model specifically designed for comprehensive image quality understanding. Unlike previous methods that depend heavily on detailed textual descriptions for supervised fine-tuning (SFT), our approach achieves superior understanding capability using only limited mean opinion scores or degradation labels.
❑ (2) We introduce a unified framework that jointly optimizes image quality rating and degradation perception, revealing mutual benefits across tasks. Within this framework, we develop three specialized rewards, including verifiable score reward, degradation classification and intensity perception rewards, enabling the GRPO framework to effectively generalize to low-level vision applications.
❑ (3) Extensive experiments across diverse datasets and IQA tasks demonstrate that Q-Insight consistently outperforms existing model-based IQA metrics as well as SFT-driven large language models. Moreover, it exhibits impressive zero-shot generalization on unseen tasks, such as reference-based image comparison reasoning, highlighting the robustness and versatility of our method.
2 Related Work
Score-based IQA methods include full-reference and no-reference approaches. Full-reference methods [47, 39, 63] assess image quality by comparing distorted images with high-quality references using traditional metrics (e.g., SSIM [47]) or advanced deep-learning-based metrics [3, 4, 9, 8, 12, 34] like LPIPS [64]. Non-reference methods evaluate quality without reference images, shifting from traditional handcrafted statistics [26, 27, 28, 29, 30, 36] to deep-learning-derived quality priors [17, 18, 23, 32, 41, 71, 72, 42, 46]. Recent multi-modal large language model (MLLM)-based methods, such as Q-Align [51] and DeQA-Score [56], leverage MLLMs’ knowledge and perceptual abilities to produce accurate scores. However, they sacrifice the intrinsic descriptive capabilities of MLLMs.
Description-based IQA methods utilize the foundational knowledge of MLLMs to deliver detailed qualitative assessments and improved interpretability [48, 50, 58, 57, 52, 5, 68, 69, 67]. For instance, Q-Bench [48] and Q-Instruct [50] enhance the low-level perceptual capabilities of MLLMs through specialized datasets and tailored evaluation strategies. Co-Instruct [52] specifically focuses on comparative quality assessments among multiple images. Approaches such as DepictQA [58] and DepictQA-Wild [57] handle both single-image and paired-image evaluations across full-reference and no-reference scenarios. Q-Ground [5] emphasizes a detailed visual quality analysis through visual grounding. However, these methods are highly dependent on extensive textual annotations for supervised fine-tuning, leading to considerable costs in human labor or GPT token consumption.
Reinforcement learning (RL) has emerged as an effective strategy to enhance the reasoning performance of LLMs through feedback-driven refinement [6, 40, 37, 53, 55, 16, 66]. Methods like RLHF [31] and RLAIF [2] employ human or AI-generated feedback to refine model behavior. In vision-language tasks, RL has successfully been employed to align model predictions closely with human preferences and reduce hallucinations [43, 59, 60, 70]. Recently, DeepSeek-R1-Zero [14] introduced Group Relative Policy Optimization (GRPO)[37], leveraging rule-based rewards to strengthen reasoning capabilities without supervised fine-tuning. Furthermore, Visual-RFT[24] applied GRPO to visual grounding, and Med-R1 [33] adopted GRPO for medical reasoning tasks. R1-VL [61] extends GRPO through step-wise optimization for multi-modal reasoning. Distinctly, our Q-Insight is the first to integrate RL-based strategies into the foundational visual quality understanding model. It jointly trains on multiple tasks and demonstrates mutually beneficial effects among them.
3 Methodology
3.1 Preliminaries
Group Relative Policy Optimization (GRPO) is an innovative reinforcement learning paradigm that has been widely used in models such as DeepSeek R1-Zero, achieving excellent results. Unlike PPO, which requires an explicit critic model to evaluate the performance of the policy model, GRPO directly computes the advantage by comparing a group of responses sampled from the policy model, greatly reducing the computational burden. Specifically, given a query $q$, GRPO samples $N$ distinct responses $\{o^{(1)},o^{(2)},\ldots,o^{(N)}\}$ from the old policy $\pi_{\theta_{\text{old}}}$. Then, the method performs the corresponding actions and receives the respective rewards $\{r^{(1)},r^{(2)},...,r^{(N)}\}$ according to the task-specific rules. By calculating the mean and standard deviation of the rewards, the relative advantages of each response can be obtained as follows:
$\hat{A}^{(i)}$ $=\frac{r^{(i)}-\operatorname{mean}(\{r^{(1)},r^{(2)}\dots,r^{(N)}\})}{\operatorname{std}(\{r^{(1)},r^{(2)}\dots,r^{(N)}\})},$ (1)
where $\hat{A}^{(i)}$ represents the normalized relative quality of the $i$-th answer. Overall, GRPO guides the policy model to prioritize higher-quality answers that receive higher reward values within the group.
After obtaining the advantage $\hat{A}^{(i)}$, GRPO calculates the ratio of the probabilities of each response under the new policy $\pi_{\theta_{\text{new}}}$ and the old policy $\pi_{\theta_{\text{old}}}$, denoted as $\text{ratio}^{(i)}$. To prevent overly large updates to the model and stabilize training, GRPO restricts the $\text{ratio}^{(i)}$ to the range $[1-\delta,1+\delta]$. To further maintain closeness to the reference distribution $\pi_{\text{ref}}$, a KL divergence penalty weighted by $\beta$ is adopted. Finally, the optimization objective of GRPO can be formulated as follows:
$\mathcal{J}(\theta)=\mathbb{E}_{[q\sim Q,o^{(i)}\sim\pi_{\theta_{\text{old}}}]}$$\left\{\min\left[\text{ratio}^{(i)},\operatorname{clip}\left(\text{ratio}^{(i)},1-\delta,1+\delta\right)\right]\hat{A}^{(i)}-\beta\cdot\mathbb{D}_{\text{KL}}[\pi_{\theta_{\text{new}}}||\pi_{\text{ref}}]\right\},$
$with~{}~{}~{}\text{ratio}^{(i)}=\pi_{\theta_{\text{new}}}(o^{(i)}~{}|~{}q)~{}/~{}{\pi_{\theta_{\text{old}}}(o^{(i)}~{}|~{}q)},$ (2)
where $Q$ denotes the candidate question set, $\mathbb{D}_{\text{KL}}$ denotes the KL regularization. $\pi_{\text{ref}}$ is typically a frozen pre-trained M-LLM. In a nutshell, GRPO effectively integrates consistent policy updates and strong reward signals in a balanced and controlled way.
3.2 Overview of Q-Insight
The overall framework of Q-Insight is illustrated in Fig. 2. During training, we jointly optimize two tasks: score regression and degradation perception. Specifically, the multi-modal input for each task comprises an image paired with a task-specific question. Given these inputs, the policy model $\pi_{\theta}$ generates groups of answers, each accompanied by explicit reasoning steps. Subsequently, each answer is evaluated using its corresponding reward function: $R_{\text{scr}}$ for score regression, and $R_{\text{deg}}$ and $R_{\text{lev}}$ for degradation perception. After computing rewards for each group of answers, the policy model is optimized jointly via the multi-task GRPO algorithm. Additionally, a KL-divergence loss is applied to constrain deviations between the policy model $\pi_{\theta}$ and the reference model $\pi_{\text{ref}}$. During inference, the trained Q-Insight generates coherent reasoning processes and outputs precise answers. Further details regarding multi-task GRPO and data construction are provided in Sec. 3.3 and Sec. 3.4.
Figure 2: Overview of the proposed Q-Insight framework. The policy model receives queries from multiple tasks and generates corresponding groups of responses accompanied by explicit reasoning steps. Task-specific reward functions ($R_{\text{scr}}$, $R_{\text{deg}}$, and $R_{\text{lev}}$) are then computed, and the policy model is subsequently optimized jointly using the multi-task group relative policy optimization algorithm.
3.3 Multi-Task Group Relative Policy Optimization
Reward design is crucial for RL-based approaches, as it allows Q-Insight to iteratively enhance its reasoning capabilities. Specifically, given each input data pair, the policy model $\pi_{\theta}$ generates a group of $N$ responses, denoted as $\{o^{(i)}\}_{i=1}^{N}$. We then evaluate these response groups using our designed reward functions to obtain $\{r^{(i)}\}_{i=1}^{N}$. This includes a general format reward shared by all tasks, along with task-specific rewards customized to the distinct characteristics of each task.
Format reward evaluates whether the reasoning steps are properly enclosed within “<think>” and “</think>” tags, and the final answer is correctly enclosed within “<answer>” and “</answer>” tags. Additionally, the content inside “<answer>” must follow a JSON-like format, beginning with “{”, ending with “}”, and containing no additional “{” or “}” characters internally. Finally, a reward score $r^{(i)}_{\text{fmt}}$ for $i$-th response of $1$ is assigned if all these conditions are satisfied. Otherwise, the reward is 00.
Rewards for score regression task. A standard way to quantify image quality is using the Mean Opinion Score (MOS). Inspired by DeepSeek-R1 [14], which leverages the final answers of math problems as a guide to promote deeper reasoning, we similarly employ MOS scores as general targets to encourage detailed reasoning and critical thinking when assessing image quality. Denoting the predicted score of the $i$-th response as $\text{scr}_{\text{pred}}^{(i)}$ and the ground-truth score as $\text{scr}_{\text{gt}}$, we design a verifiable score reward function $R_{\text{scr}}$ as follows. The reward value $r_{\text{scr}}^{(i)}$ for the $i$-th response is determined by:
$r^{(i)}_{\text{scr}}=1\ \quad\text{if}\ \mid\!\text{scr}^{(i)}_{\text{pred}}-\text{scr}_{\text{gt}}\!\mid<\epsilon,~{}\text{otherwise }0,$ (3)
where $\epsilon$ is a predefined threshold, allowing the model’s score predictions to fluctuate within an acceptable range rather than requiring exact accuracy. Thus, the predicted score receives a reward of $1$ if it falls within the threshold $\epsilon$ of the ground-truth MOS, and 00 otherwise.
Table 1: Prompts for Different Tasks. The system prompt is shared across all tasks, while task-specific prompts are additionally designed for each individual task.
System Prompt: A conversation between User and Assistant. The user asks a question, and the Assistant solves it. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> reasoning process here </think><answer> answer here </answer>.
Prompt for Score Regression Task: What is your overall rating on the quality of this picture? The rating should be a float between 1 and 5, rounded to two decimal places, with 1 representing very poor quality and 5 representing excellent quality. Return the final answer in JSON format with the following keys: "rating": The score.
Prompt for Degradation Perception Task: Analyze the given image and determine if it contains any of the following distortions: "noise", "compression", "blur", or "darken". If a distortion is present, classify its severity as "slight", "moderate", "obvious", "serious", or "catastrophic". Return the result in JSON format with the following keys: "distortion_class": The detected distortion (or "null" if none). and "severity": The severity level (or "null" if none).
Prompt for Reference-Based Zero-shot Comparison Reasoning: Given reference low-quality image: <ref_image>, Image A: <image_A> and Image B: <image_B>. Image A and Image B are results from two resolution methods. Explain and justify whether Image A or Image B is better. You should think from the following aspects: (1) Content fidelity and consistency with the reference image and (2) overall quality. Your answer should be "Image A" or "Image B".
Rewards for degradation perception task. In the degradation perception task, the model is expected to predict both the distortion class and its corresponding level. Denoting the predicted distortion class and level of the $i$-th response as $\text{deg}^{(i)}_{\text{pred}}$ and $\text{lev}^{(i)}_{\text{pred}}$ respectively, we define the degradation classification reward as follows. The reward value $r^{(i)}_{\text{deg}}$ is determined by:
$r^{(i)}_{\text{deg}}=1\quad\text{if}\ \text{deg}^{(i)}_{\text{pred}}=\text{deg}_{\text{gt}},~{}\text{otherwise }0.$ (4)
This means the predicted distortion class receives a reward of $1$ if correct, and 00 otherwise. Similarly, the distortion-level reward $r^{(i)}_{\text{lev}}$ is determined by:
$r^{(i)}_{\text{lev}}=1\quad\text{if}\ \text{deg}^{(i)}_{\text{pred}}=\text{deg}_{\text{gt}}\ {\text{and}}\ \text{lev}^{(i)}_{\text{pred}}=\text{lev}_{\text{gt}},~{}\text{otherwise }0.$ (5)
Thus, the predicted distortion level earns a reward of $1$ only if both the predicted class and the level exactly match the ground truth; otherwise, it receives 00.
Finally, the overall reward of $i$-th response is calculated as:
$r^{(i)}=r^{(i)}_{\text{fmt}}+\mathds{1}_{\text{scr}}\cdot r^{(i)}_{\text{scr}}+\mathds{1}_{\text{deg}}\cdot\left(\alpha_{1}\cdot r^{(i)}_{\text{deg}}+\alpha_{2}\cdot r^{(i)}_{\text{lev}}\right),$ (6)
where $\mathds{1}_{\text{scr}}$ equals $1$ if the score regression task is selected (and 00 otherwise), and similarly, $\mathds{1}_{\text{deg}}$ equals $1$ if the degradation perception task is selected (and 00 otherwise). After computing rewards for all generated responses $\{r^{(1)},r^{(2)},...,r^{(N)}\}$, the policy model is updated following Eqs. (1) and (3.1).
3.4 Data Construction
We construct multi-modal training data to jointly train Q-Insight on the score regression and degradation perception tasks. The prompts designed for each task are detailed in Tab. 1. Specifically, a general system prompt is shared across tasks, which encourages the model to explicitly output its reasoning process and provide structured responses. This general prompt is supplemented by task-specific prompts tailored for score regression and degradation perception, respectively. For the score regression task, the input includes a task-specific prompt and the image to be rated, with the Mean Opinion Score (MOS) serving as the ground-truth. In the degradation perception task, the input consists of a prompt and an image characterized by a specific distortion class and severity level. We define five distortion categories: “noise”, “blur”, “JPEG”, “darken”, and “null”, where “null” indicates no distortion. Each distortion type has five severity levels: “slight”, “moderate”, “obvious”, “serious”, and “catastrophic”. The distortion class and corresponding severity level constitute the ground-truth labels. Additionally, for the zero-shot comparative reasoning scenario, inputs include a prompt, two images to be compared, and an optional reference image.
Table 2: PLCC / SRCC comparison on the score regression tasks between our Q-Insight and other competitive IQA methods. All methods except handcrafted ones are trained on the KonIQ dataset. Our Q-Insight outperforms all baseline methods across nearly all benchmarks.
{NiceTabular}
[…]
Category Methods KonIQ SPAQ KADID PIPAL LiveW AGIQA CSIQ AVG.
Handcrafted NIQE [28] 0.533 0.679 0.468 0.195 0.493 0.560 0.718 0.521
(SPL 2012) / 0.530 / 0.664 / 0.405 / 0.161 / 0.449 / 0.533 / 0.628 / 0.481
BRISQUE [27] 0.225 0.490 0.429 0.267 0.361 0.541 0.740 0.436
(TIP 2012) / 0.226 / 0.406 / 0.356 / 0.232 / 0.313 / 0.497 / 0.556 / 0.369
Non-MLLM Deep-learning NIMA [44] 0.896 0.838 0.532 0.390 0.814 0.715 0.695 0.697
(TIP 2018) / 0.859 / 0.856 / 0.535 / 0.399 / 0.771 / 0.654 / 0.649 / 0.675
HyperIQA [41] 0.917 0.791 0.506 0.410 0.772 0.702 0.752 0.693
(CVPR 2020) / 0.906 / 0.788 / 0.468 / 0.403 / 0.749 / 0.640 / 0.717 / 0.667
DBCNN [65] 0.884 0.812 0.497 0.384 0.773 0.730 0.586 0.667
(TCSVT 2020) / 0.875 / 0.806 / 0.484 / 0.381 / 0.755 / 0.641 / 0.572 / 0.645
MUSIQ [18] 0.924 0.868 0.575 0.431 0.789 0.722 0.771 0.726
(ICCV 2021) / 0.929 / 0.863 / 0.556 / 0.431 / 0.830 / 0.630 / 0.710 / 0.707
CLIP-IQA+ [46] 0.909 0.866 0.653 0.427 0.832 0.736 0.772 0.742
(AAAI 2023) / 0.895 / 0.864 / 0.654 / 0.419 / 0.805 / 0.685 / 0.719 / 0.720
ManIQA [54] 0.849 0.768 0.499 0.457 0.849 0.723 0.623 0.681
(CVPR 2022) / 0.834 / 0.758 / 0.465 / 0.452 / 0.832 / 0.636 / 0.627 / 0.658
MLLM-based C2Score [73] 0.923 0.867 0.500 0.354 0.786 0.777 0.735 0.706
(NeurIPS 2024) / 0.910 / 0.860 / 0.453 / 0.342 / 0.772 / 0.671 / 0.705 / 0.673
Qwen-SFT [1] 0.889 0.874 0.668 0.473 0.734 0.813 0.674 0.732
(Arxiv 2025) / 0.866 / 0.875 / 0.663 / 0.442 / 0.728 / 0.739 / 0.650 / 0.709
Q-Align [51] 0.941 0.886 0.674 0.403 0.853 0.772 0.671 0.705
(ICML 2024) / 0.940 / 0.887 / 0.684 / 0.419 / 0.860 / 0.735 / 0.737 / 0.752
DeQA [56] 0.953 0.895 0.694 0.472 0.892 0.809 0.787 0.786
(CVPR 2025) / 0.941 / 0.896 / 0.687 / 0.478 / 0.879 / 0.729 / 0.744 / 0.765
Q-Insight 0.933 0.907 0.742 0.486 0.893 0.811 0.870 0.806
(Ours) / 0.916 / 0.905 / 0.736 / 0.474 / 0.865 / 0.764 / 0.824 / 0.783
4 Experiments
4.1 Experimental Setup
Datasets and Metrics. For the score regression task, we use diverse IQA datasets across four categories: (a) in-the-wild datasets, including KonIQ [15], SPAQ [10], and LIVE-Wild [11]; (b) synthetic distortion datasets, including KADID [21] and CSIQ [19]; (c) model-processed distortions, including PIPAL [13]; and (d) AI-generated images from AGIQA [20]. Following [56], we split KonIQ into training and test sets, with approximately $7000$ training images. Mean Opinion Scores (MOS) across these datasets are normalized into the range $[1,5]$. The remaining datasets are exclusively used to evaluate the model’s out-of-distribution (OOD) generalization capability. For the degradation perception task, we randomly select $7000$ images from DQ-495K [57] that contain a single distortion for training, with an additional $1000$ images reserved for testing. We adopt the pearson linear correlation coefficient (PLCC) and spearman rank-order correlation coefficient (SRCC) as metrics to evaluate performance on the score regression task, following [18, 51, 56]. For the degradation perception task, we employ distortion class accuracy and degradation level accuracy as evaluation metrics.
Implementation Details. We adopt Qwen-2.5-VL-7B [1] as our base model. In the GRPO algorithm, the generation number $N$ is set to 8, while the weights $\alpha_{1}$ and $\alpha_{2}$ are set to $0.25$ and $0.75$, respectively. The threshold $\epsilon$ is set to $0.35$. We employ AdamW [25] as the optimizer, using an initial learning rate of $1\times 10^{-6}$ that linearly decays during training. The model is trained for 10 epochs with a total batch size of 128. Training is completed in approximately one day using 16 NVIDIA A100 GPUs.
Figure 3: Score rating and explanation results of our Q-Insight. Q-Insight is capable of recognizing text, analyzing the lighting and shading conditions of an image, and understanding its composition.
Table 3: Distortion prediction accuracy (Deg. Acc.) and degradation level accuracy (Lev. Acc.) comparison between our Q-Insight and AgenticIR [74]. Our method outperforms AgenticIR across all degradations, especially in Noise and JPEG Compression.
{NiceTabular}
[…]
AgenticIR [74] (ICLR 2025) Deg. Acc. 0.4646 0.8390 0.0135 0.7478 0.9339 0.5998
Lev. Acc. 0.1858 0.3219 0.0000 0.2611 - 0.1922
Q-Insight (Ours) Deg. Acc. 1.0000 0.9756 1.0000 0.9027 0.7603 0.9277
Lev. Acc. 0.5973 0.4438 0.5541 0.3230 - 0.4796
4.2 Score Regression
We first evaluate our Q-Insight on the score regression task. We compare Q-Insight with handcrafted methods NIQE [28] and BRISQUE [27]; non-MLLM deep-learning methods including NIMA [44], HyperIQA [41], DBCNN [65], MUSIQ [18], CLIP-IQA+ [46], and ManIQA [54]; and recent MLLM-based methods such as C2Score [73], Q-Align [51], DeQA-Score [56], and a supervised fine-tuned Qwen [1]. For a fair comparison, all methods (except handcrafted ones) are trained on the KonIQ dataset, and all MLLM-based methods utilize approximately 7B parameters. The comparison results in terms of PLCC and SRCC between Q-Insight and other IQA methods are presented in Tab. 3.4. Compared with the state-of-the-art method DeQA-Score, our Q-Insight performs slightly worse on the in-domain KonIQ dataset. However, on out-of-distribution (OOD) datasets, Q-Insight consistently outperforms all baseline methods across nearly all benchmarks, achieving approximately $0.2$ improvements in both PLCC and SRCC. This demonstrates the effectiveness and strong generalization capability of our approach. Fig. 3 shows two illustrative cases of the reasoning process in the score regression task. Notably, our Q-Insight does not merely output a numerical score; it also provides a comprehensive and detailed reasoning process. In the first example (top of Fig. 3), Q-Insight successfully recognizes the text on the neon sign and carefully analyzes details such as lighting conditions. Similarly, the second example (bottom of Fig. 3) highlights Q-Insight’s capability in understanding image composition. This further demonstrates that Q-Insight moves beyond merely regressing scores, critically analyzing image quality issues from multiple perspectives, and advancing toward a more comprehensive understanding of image quality.
Table 4: Ablation study on the score regression task between multi-task and single-task training. Ours with joint-training significantly outperforms score-only training on PLCC / SRCC metrics.
{NiceTabular}
[…]
Ours 0.918 0.903 0.702 0.458 0.870 0.816 0.685 0.765
(Score-Only) / 0.895 / 0.899 / 0.702 / 0.435 / 0.839 / 0.766 / 0.640 / 0.739
Ours 0.933 0.907 0.742 0.486 0.893 0.811 0.870 0.806
(Joint-Training) / 0.916 / 0.905 / 0.736 / 0.474 / 0.865 / 0.764 / 0.824 / 0.783
Table 5: Ablation study on the degradation perception task between multi-task and single-task training. Jointly training with score regression improves the accuracy of degradation perception.
{NiceTabular}
[…]
Ours (Degradation-Only) Deg. Acc. 0.9867 0.9268 0.9685 0.8805 0.5702 0.8960
Lev. Acc. 0.4343 0.3951 0.3108 0.2567 - 0.3492
Ours (Joint-Training) Deg. Acc. 1.0000 0.9756 1.0000 0.9027 0.7603 0.9277
Lev. Acc. 0.5973 0.4438 0.5541 0.3230 - 0.4796
4.3 Distortion Perception
We further evaluate Q-Insight on the distortion perception task, comparing it with AgenticIR [74], which fine-tunes an MLLM to perform a similar distortion prediction function. The comparative results are presented in Tab. 4.1. Notably, AgenticIR requires sequential queries for each possible distortion type, whereas Q-Insight identifies distortion types using only a single query. Q-Insight consistently outperforms AgenticIR across nearly all distortion categories, resulting in significantly higher average accuracy. However, for the “null” category (no distortion), our performance is slightly lower than AgenticIR, possibly because the reasoning mechanism of Q-Insight is more sensitive to the presence of potential distortions. Overall, these results highlight Q-Insight’s ability to efficiently and accurately perceive image distortions through a concise and unified reasoning framework.
Figure 4: Subjective ablation comparison between joint multi-task training and w/o joint training on the explanation of image scoring. With joint training, our method can better perceive degradation cues in images (such as pixelated appearance), thereby improving the accuracy of quality assessment.
Figure 5: Image comparison reasoning results of our Q-Insight and DepictQA [58, 57]. Q-Insight outperforms DepictQA in comprehensive content understanding and accurate degradation perception.
4.4 Ablation Studies
Effect of multi-task training. To demonstrate the effectiveness of multi-task training, we compare our jointly trained Q-Insight with two single-task variants, trained separately on each task. The comparison results are presented in Tabs. 4.2 and 4.2. As shown in Tab. 4.2, the jointly trained Q-Insight significantly outperforms the score-only variant on nearly all datasets, particularly those with synthetic distortions (KADID, CSIQ) and model-processed distortions (PIPAL). This clearly indicates that incorporating the degradation perception task benefits the score regression task. Fig. 4 further illustrates the importance of multi-task training: Q-Insight can better perceive detailed degradations such as pixel-level artifacts, thus enhancing overall accuracy in quality assessment. Similarly, Tab. 4.2 demonstrates that in the degradation perception task, the jointly trained model outperforms the degradation-only variant across all distortion categories. This suggests that the score regression task also positively contributes to the degradation perception capability. These results confirm the mutual benefit and effectiveness of our multi-task training approach. Furthermore, this clearly demonstrates that the visual quality understanding potential of MLLMs can be significantly enhanced through carefully designed training tasks and objectives.
4.5 Zero-Shot Image Comparison Reasoning
Our Q-Insight effectively generalizes to zero-shot image comparison reasoning tasks, covering both reference-based and non-reference-based scenarios, as illustrated in Figs. 1 and 5. In Fig. 1, Q-Insight conducts a reference-based comparison, where the reference image is of lower quality, and Images A and B are outputs from different super-resolution methods. Q-Insight compares these two images from both fidelity and overall quality perspectives, demonstrating its potential utility in image enhancement tasks. In Fig. 5, Q-Insight surpasses DepictQA [58, 57] in terms of comprehensive content understanding and accurate degradation perception. Importantly, Q-Insight achieves this without relying on specialized training on large-scale textual datasets, highlighting its strong generalization ability enabled by the GRPO-based framework and multi-task training strategy.
5 Conclusion
In this paper, we introduce Q-Insight, a novel GRPO-based model for comprehensive image quality understanding. It jointly optimizes score regression and degradation perception tasks using only a limited amount of labeled data. Unlike traditional methods that rely on extensive textual annotations or purely numerical scoring, our framework combines numerical accuracy with interpretative reasoning, significantly improving the perceptual analysis capabilities of image quality models. Extensive experiments show that Q-Insight consistently outperforms existing state-of-the-art methods across various datasets and tasks, demonstrating impressive zero-shot generalization to image comparison reasoning. Looking ahead, Q-Insight can extend its capabilities to a wide range of tasks, such as image aesthetic assessments, and serve as a powerful discriminative signal to improve image enhancement models. As a unified model for scoring, perception, comparison, and reasoning, Q-Insight can act as a central hub, coordinating image reconstruction tools and providing valuable insights into the enhancement process. This integrated and automated system has the potential to revolutionize image quality understanding and enhancement, providing a unified solution that can transform how image quality is evaluated, improved, and applied across various fields.